OpenStudy (anonymous):

What does x(x^0.5) = ?

OpenStudy (anonymous):

If x^(3/2) = 125000, what does x equal?

OpenStudy (jdoe0001):

x^(a/b) = b root of x^a

OpenStudy (anonymous):

is it not a root of x^b?

OpenStudy (jdoe0001):

a/b, b would be the root, a the exponent

OpenStudy (anonymous):

\[x(x^{0.5})=x^1\times x^{0.5}=x^{1.5}=x^{3/2}\\ x^{3/2}=125000\\ \text{lets take cube root}\\ (x^{3/2})^{1/3}=(125000)^{1/3}\\ x^{1/2}=\ldots\\ \text{now square both sides}\\ x=\ldots. \]